### v1

# TartanGround: A Large-Scale Dataset for Ground Robot Perception and Navigation

Abstract
We present TartanGround, a large-scale, multi-modal dataset to advance the perception and autonomy of ground robots operating in diverse environments. This dataset, collected in various photorealistic simulation environments includes multiple RGB stereo cameras for 360-degree coverage, along with depth, optical flow, stereo disparity, LiDAR point clouds, ground truth poses, semantic segmented images, and occupancy maps with semantic labels. Data is collected using an integrated automatic pipeline, which generates trajectories mimicking the motion patterns of various ground robot platforms, including wheeled and legged robots. We collect 910 trajectories across 70 environments, resulting in 1.5 million samples. Evaluations on occupancy prediction and SLAM tasks reveal that state-of-the-art methods trained on existing datasets struggle to generalize across diverse scenes. TartanGround can serve as a testbed for training and evaluation of a broad range of learning-based tasks, including occupancy prediction, SLAM, neural scene representation, perception-based navigation, and more, enabling advancements in robotic perception and autonomy towards achieving robust models generalizable to more diverse scenarios. The dataset and codebase for data collection will be made publicly available upon acceptance.
Webpage: https://tartanair.org/tartanground
I Introduction
Ground robots are increasingly being used in a wide range of environments, from structured urban areas to unstructured terrains such as forests, farmlands, and construction sites. These robots serve various purposes, including autonomous delivery, agricultural automation, industrial inspection, search-and-rescue missions, and construction site monitoring. To improve their adaptability and generalizability in these diverse settings, data-driven methods have gained traction. These approaches tackle key tasks in perception and scene understanding, such as , occupancy prediction, semantic segmentation, monocular depth estimation, etc, enabling robots to better interpret, interact, and navigate their surroundings.
In the domain of autonomous driving, large-scale datasets [1, 2, 3, 4] have played a crucial role in advancing machine learning models for tasks such as object detection, occupancy prediction, and semantic segmentation. These datasets offer standardized benchmarks that allow for consistent evaluation and comparison of algorithms. However, there is a notable lack of similar datasets and benchmarks for mobile robots operating in a broader range of environments. This absence makes it challenging to develop and evaluate generalizable models that can perform reliably across various and complex settings.
Several datasets support research in mobile robotics, each tailored to specific environments. For example, datasets [5, 6, 7, 8, 9, 10] collected in off-road environments provide sensor modalities ranging from RGB cameras and LiDARs for scene understanding, to proprioceptive and traction data for vehicle dynamics modeling. Indoor datasets such as ScanNet [11], TUM RGB-D [12], and Matterport3D [13] provide RGB-D data for 3D reconstruction and . However, these datasets often have limitations in environmental diversity, size, accurate ground truths, or sensor types, which limit their applicability in developing robust, generalizable models.
Simulation datasets have become an important alternative to overcome these difficulties and limitations of real-world data collection in robotics. They offer a controlled environment for generating large-scale, high-quality data with precise ground truth annotations for semantics, depth, and poses—information that is often difficult to obtain in real-world settings. Furthermore, simulation enables diverse data collection under varying lighting, weather, and terrain conditions, improving model robustness and generalization. These benefits have made simulation datasets crucial for advancing perception tasks such as and monocular depth estimation. For instance, TartanAir [14] has been instrumental in training algorithms like DROID-SLAM [15], demonstrating effective sim-to-real transfer. Similarly, synthetic datasets [16, 17, 18] have played a key role in training foundation models for monocular depth estimation [19, 20], further showcasing the impact of learning from simulation data.
To address the need for comprehensive data resources, we introduce TartanGround, a large-scale simulation dataset designed to support perception and navigation tasks for ground robots across diverse environments. TartanGround includes data from over 70 diverse, challenging photorealistic environments, offering multi-modal sensor data such as multiple stereo RGB-D images covering a 360°, semantic labels, LiDAR point clouds, semantic occupancy maps, and ground truth poses. We show that the existing occupancy prediction methods, solely trained on data from the autonomous driving domain, do not generalize to environments of mobile robot operations such as forests. Moreover, we evaluate various algorithms and find that they struggle in challenging scenarios of low visibility and heavy occlusions. To summarize, the main contributions of this paper are: (1) A large dataset with over 1.5 million samples collected across diverse environments with precise ground truth labels mimicking different ground robot motion patterns, including wheeled (omnidirectional and differential-drive), and legged robots (quadrupedal), (2) an automatic data collection pipeline and (3) evaluation of two tasks, occupancy prediction and highlighting the limitations of models trained on existing datasets.
II Related Work
In the field of autonomous driving, the KITTI [1] dataset offers a comprehensive suite of sensor data, including stereo camera and LiDAR inputs, primarily for autonomous driving research. Building upon this, SemanticKITTI [21] extends KITTI by providing dense point-wise semantic labels for LiDAR scans, enabling advancements in LiDAR segmentation tasks. NuScenes [4] and Waymo Open Dataset [2] further contribute to urban scene understanding by offering 360 sensor coverage and diverse urban scenarios. Occ3D [22] establishes new benchmarks for semantic occupancy prediction on nuScenes and Waymo, facilitating the development of models that predict both the geometry and semantics of urban environments.
In indoor environments, several RGB-D datasets have been instrumental in advancing scene understanding. ScanNet [11] comprises annotated 3D reconstructions of indoor scenes, serving as a benchmark for tasks like 3D semantic segmentation and object recognition. TUM RGB-D [12] offers sequences recorded with handheld cameras, providing ground truth trajectories for evaluating SLAM and odometry algorithms. SUN RGB-D [23] and NYUv2 [24] provide paired RGB-D with semantic labels, supporting research in semantic segmentation and monocular depth estimation. Matterport3D [13] offers a rich collection of indoor images, facilitating research in 3D reconstruction and navigation.
For off-road environments, RELLIS-3D [5] provides multi-modal data with dense annotations, supporting semantic segmentation research. RUGD [6] offers images captured in unstructured outdoor environments with pixel-wise semantic labels. GOOSE [7] presents data collected with a vehicle equipped with multiple cameras and LiDARs ensuring 360 coverage in offroad environments. WildScenes [10] provides synchronized image and LiDAR data with semantic annotations in natural settings, while TartanDrive [8] offers extensive sensor data for learning dynamics models in off-road driving scenarios. WildOcc [25] builds upon RELLIS-3D by providing semantic occupancy annotations, enabling research in 3D scene understanding in off-road context.
Despite the significant contributions of these datasets, they are often limited by scale, environmental diversity, or sensor modalities, which can hinder the development of generalizable models for robotic perception and navigation. To address these limitations, TartanAir [14] introduced a large-scale synthetic dataset with 20 diverse environments and random motion patterns, enhancing generalization by offering photo-realistic imagery under varied weather and lighting conditions. TartanAir-V2 extended V1 with more scenes and more modalities [26]. Building on this, we introduce TartanGround, specifically targeting ground robots. It features realistic ground robot motion patterns, including wheeled and legged robots. We summarize the various datasets along with the sensor data, available ground truth, and scale in Tab. I. TartanGround is the only large-scale dataset covering diverse environments and having realistic ground robot motions.
| Dataset | Env | Motion Pattern | Stereo | Depth | LiDAR | Semantics | Occupancy | Pose | Multi-cam | Seq Num | Samples |
| Semantic KITTI [21] | Urban | Car | ✔ | ✘ | ✔ | ✔ | ✔ | ✔ | ✘ | 22 | 43 K |
| Occ3D-Waymo [2, 22] | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 5 | 1000 | 200 K |
| Occ3D-nuScenes [4, 22] | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 6 | 1000 | 40 K |
| RELLIS-3D [5] | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 6.2 K |
| RUGD [6] | Off-road | Car | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | 18 | 7.5 K |
| TartanDrive [8] | Off-road | Car | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | 630 | 184 K |
| WildScenes [10] | Off-road | Hand | ✘ | ✘ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 9.6 K |
| WildOcc [25] | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✘ | 5 | 10 K |
| GOOSE [7] | Off-road | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 356 | 10 K |
| Scannet [11] | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✔ | ✔ | ✘ | 1500 | 2.5 M |
| SUN RGBD [23] | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | - | 10 K |
| GrandTour [27] | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✘ | ✘ | ✔ | 3 | 100 | 100 K |
| TartanAir [14] | Mix | Random | ✔ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | 1037 | 400 k |
| TartanAir-V2 [26] | Mix | Random | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 1110 | 1.4 M |
| TartanGround | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 910 | 1.5 M |
III The Dataset
III-A Features
TartanGround consists of 70 realistic simulation environments from TartanAir-V2 [26], covering diverse scenarios from structured urban to large-scale unstructured outdoor environments (Fig. 2). The environments have been developed with Unreal Engine, and the data is collected using the AirSim [28] plugin. This setup allows for rendering photorealistic scenes with high fidelity and makes it possible to have dynamic lighting, adverse weather effects, dynamic objects, and seasonal changes. In each of the environments, we collect data from multiple sampled trajectories with diverse motion patterns mimicking real-world ground robots. To achieve full 360 coverage, we use 6 stereo RGB cameras (front, left, right, back, top, bottom), each with a of 90. Each camera is synchronized with accurate ground truth poses and also records depth and semantic segmentation images. During post-processing, we generate additional ground truth data, including IMU data, optical flow, stereo disparity, LiDAR point clouds, and semantic occupancy maps. In total, we collect 910 trajectories (441 omni-wheeled, 226 diff-wheeled, and 243 legged), each with 600 to 8000 samples, resulting in a total dataset size of approximately 15 TB, 1.5 million samples, and 18 million RGB images. This large-scale, multi-modal dataset is designed to support a wide range of robotic perception and navigation tasks, benefiting the community in establishing new benchmarks for generalizable learning-based methods.
III-B Trajectory Sampling
The overview of the trajectory sampling pipeline is shown in Fig. 3.
III-B1 Environment Pointcloud
We export the corresponding mesh (Fig. 3a) from Unreal Engine for each environment and sample a point cloud (Fig. 3b) with a specified density . We exclude foliage elements such as grass and small bushes from the mesh export to ensure our geometry-based traversability estimation accurately reflects navigable paths. This approach allows us to generate trajectories that traverse compressible vegetation, capturing more realistic and diverse data.
III-B2 Traversability Generation
We utilize a geometry-based traversability estimation pipeline inspired by [29], which, taking as an input the environment point cloud, can efficiently represent the environment with complex terrain conditions and spatial structures into 3D tomogram slices (Fig. 3c). This method extends traditional elevation maps [30] for multi-layered environments by assigning ground and ceiling elevations at fixed intervals, facilitating efficient planning of 3D trajectories for ground robots. Moreover, it incorporates the robot’s capabilities, such as climbing stairs and navigating steep gradients, by assessing factors like slope, step height, and overhead clearance to compute the traversability values. This approach ensures robust performance in both structured and unstructured multi-layered environments.
III-B3 Sparse Trajectory Sampling
Our objective is to generate long trajectories that maximize coverage of the environment while maintaining spatial diversity (Fig. 3d). To achieve this, we begin by uniformly sampling points in free space and points near obstacles, ensuring that the trajectories encompass both open and constrained regions. To eliminate redundancy, we apply a representative point sampling step using k-means clustering, resulting in a set of representative points.
Each representative point is randomly assigned to one of the trajectory subgroups. For each subgroup, we construct a graph where nodes correspond to the sampled points and edges represent the path distance between them in the tomogram. We approximate the Traveling Salesman Problem (TSP) on this graph to determine an optimal traversal sequence. For each pair of consecutive nodes in the traversal sequence, we find the optimal smooth path using an A approach for tomograms [29]. The final path is obtained by concatenating the subpaths between consecutive nodes, ensuring connectivity. This approach is summarized in Alg. 1. By the end of this process, we obtain sparse trajectories that provide comprehensive coverage of the environment.
III-B4 Dense Trajectory Generation
To collect data in AirSim, we generate dense poses at a fixed frequency of , requiring realistic interpolation of sparse trajectories to accurately mimic ground robot motion while adhering to velocity and acceleration constraints. We implement three trajectory variations: (1) omnidirectional, (2) differential-drive (Fig. 3e), and (3) legged robot trajectories (Sec. III-C).
We use a fixed look-ahead distance for path tracking, ensuring the robot moves toward its tracking point. In omnidirectional motion, the robot can translate in any direction on the ground plane, meaning its heading may not always align with its movement direction. In contrast, the differential-drive model enforces a stricter motion constraint, requiring the robot to reorient itself to the next tracking point before proceeding. We apply a random walk model to introduce realistic velocity variations, adjusting speeds dynamically while sampling acceleration values within predefined limits at each timestep. Additionally, we ensure smooth yaw transitions using a bounded yaw rate, maintaining physically plausible motion. We randomly sample the robot height in the range for each trajectory. We also introduce Gaussian noise to the position to simulate real-world uncertainties, adding slight perturbations that account for terrain roughness and sensor and actuation inconsistencies.
III-C Legged Robot Trajectories
To capture realistic motion patterns of a legged robot, we perform path tracking of the sparse trajectories within a Gazebo simulation environment using an ANYmal D legged robot (Fig. 3f). An example of the ANYmal robot in action in a forest environment is shown in Fig. 1.
During simulation, we record the ground truth base poses, which are later used for sampling photorealistic data in AirSim. In addition to the base poses, we collect a comprehensive set of proprioceptive data, including base velocities and accelerations, joint states, and contact forces for each leg. In addition to the perception tasks, these trajectories can also be used for learning navigation tasks.
III-D Data Collection, Verification and Post-processing
For the provided dense poses (Fig. 3e-f), we capture 6 RGB stereo image pairs in AirSim, along with corresponding depth and semantic segmentation images. Using an approach similar to TartanAir [14], we generate additional ground truth data, including optical flow, stereo disparity, and simulated LiDAR measurements from these raw images. To enhance the usability of the dataset, we introduce a custom camera resampling feature that enables image extraction with arbitrary intrinsics and rotation matrices. This allows users to specify camera parameters that match their real robot setup, and the system re-renders images accordingly from the captured set of 6 images, ensuring compatibility with diverse robotic platforms.
For data verification, we ensure the synchronization between camera poses and captured images by computing the optical flow between consecutive image pairs and evaluating the mean photometric error, as described in TartanAir [14]. Additionally, depth images are analyzed to verify collision occurrences with the environment.
IV Experiments and Applications
In this section, we evaluate the performance of state-of-the-art methods on two key tasks: Occupancy Prediction and SLAM. We further discuss the other potential applications of the dataset.
IV-A Occupancy Prediction
The task of predicting 3D occupancy voxels using multi-camera images has become quite popular in the field of autonomous driving as it enables detailed spatial representation of the environment useful for downstream tasks such as path planning and navigation. End-to-end occupancy prediction networks [22] have the advantage of handling occlusions and satisfying multi-camera consistency where traditional methods struggle. NuScenes [4] and Waymo [2] have become popular benchmarks for evaluating these learned networks, however, there is a lack of similar benchmarks for mobile robots operating in diverse environments. We show through experiments that the methods trained on autonomous driving data do not generalize to other environments, and thus, there is a need for such a large-scale dataset for training and evaluating in different scenes.
IV-A1 Setup, Baselines, and Environments
We set up two baselines for occupancy prediction. The first is a simple baseline that uses an off-the-shelf monocular depth estimator [20] to project pixels into 3D space using the predicted depth. To reduce the effect of bleeding artifacts, we further apply gradient filtering. This baseline is designed to highlight the limitations of depth-based projections and emphasize the need for dedicated occupancy prediction networks. Our second baseline is SurroundOcc [31], a state-of-the-art 3D occupancy prediction network. This network takes multiple RGB images as input and extracts per-image multi-scale features. Using 2D-3D spatial attention, the multi-camera information is fused to construct 3D multi-scale feature volumes which are decoded into semantic occupancy predictions.
For evaluating the baselines, we select three trajectories each from urban and natural environments, and truncate them to 1000 samples per trajectory. The urban environments depict city-like scenarios that are closer to the training domain of scenes (Fig. 4), while the natural environments include forest environments in different seasons and a marsh environment with heavy fog and low visibility. SurroundOcc network was trained on nuScenes data which was collected using a car mounted with six cameras having overlapping and facing front-left, front, front-right, back-left, back and back-right directions at a resolution of 1600x900 pixels. To minimize the domain gap, we re-render images matching this setup using our image resampling pipeline.
We use the metric to evaluate the performance of occupancy prediction. In principle, the network also predicts the semantic class along with the occupancy, however, since the labels of nuScenes do not match our environment labels, we do not evaluate this. Moreover, instead of evaluating in the range of from the egocentric frame in x-y directions as in SurroundOcc, we only evaluate in the range of (at resolution) due to the relatively lower height of the cameras in our case limiting the longer distance visibility.
IV-A2 Evaluations
The quantitative results are summarized in Tab. II. In general, we observe that SurroundOcc outperforms the depth-based projection pipeline across all urban environments. This is expected since urban environments have a similar distribution as the nuScenes training dataset. Qualitative results for the ModNeighborhood environment are shown in Fig. 5. Here we see that SurroundOcc is able to predict the occupancy as well as the semantic classes (which we do not evaluate) for the car, vegetation, and driveable surface. On the other hand, visualizing the results of the depth projection method, we can clearly see the effect of bleeding and the inability to handle occlusions. Moreover, we also observe that this method suffers from inconsistent predictions across multiple views, leading to duplicating of objects when projected into 3D. These limitations highlight the advantages of occupancy prediction networks.
The natural environments are completely out-of-distribution for SurroundOcc, and thus, the performance in these environments is significantly lower than the urban environments. Moreover, these environments are also more challenging, with lower visibility and higher occlusion due to the presence of dense trees and tall grass. Interestingly, here, the depth projection method performs much better highlighting the generalization capabilities of the monocular depth-estimator network. We believe that our large-scale dataset can contribute towards the development of robust generalizable models for predicting sematic occupancy in diverse environments.
IV-B Visual Odometry and SLAM
Visual Odometry (VO) and SLAM for ground robots pose unique challenges, such as vegetation occlusions, short horizon caused by low altitude, and aggressive motion due to bumpy terrain. TartanGround is designed to capture these challenging cases (as shown in Fig. 6). As a result, it becomes an interesting benchmark for VO and SLAM research.
IV-B1 Testing Environments and Baselines
In this section, we show a small-scale evaluation of three state-of-the-art VO/SLAM algorithms on four challenging trajectories from four different environments. The trajectory from the Forest environment contains occlusion from tall grass and bushes, which are commonly presented in real-world off-road scenes. CastleFortress is a large scene with indoor-outdoor transitions, bringing dramatic illumination change. WaterMillDay has a rushing creek with water splash. ModUrbanCity contains a narrow dark stair that lacks good visual features. Our baseline consists of a classic geometry-based SLAM algorithm ORB-SLAM3 [33], a learning-based monocular odometry DPVO [34], and a stereo VO model MACVO [35]. We use relative translation error (, ) and relative rotation error (, ) to evaluate the results.
IV-B2 Evaluation
Table III presents the quantitative results of various VO/SLAM algorithms on the TartanGround trajectories. To ensure a fair comparison, we interpolate missing data points and concatenate successfully tracked segments when algorithms lose tracking. Among the evaluated algorithms, ORB-SLAM3 frequently loses tracking, leading to significantly elevated translation and rotation errors. In contrast, MACVO and DPVO exhibit robustness across most environments. Notably, DPVO’s multi-frame design makes it more accurate in orientation estimation. However, its performance degrades in the presence of visual occlusions. MACVO is a frame-to-frame stereo odometry. It suffers from sudden viewpoint shifts but still shows exceptional translation accuracy. TartanGround poses a unique challenge for the existing SLAM algorithms, and becomes a good complement to the existing SLAM benchmarks.
| Scenes | ORB-SLAM3 [33] | DPVO⋆† [34] | MAC-VO [35] | |||
| Forest | 0.152 | 1.667 | 0.010 | 0.039 | 0.008 | 0.038 |
| CastleFortress | 0.338 | 1.943 | 0.031 | 0.041 | 0.014 | 0.070 |
| WaterMill | 0.404 | 1.090 | 0.089 | 0.036 | 0.016 | 0.062 |
| ModUrban | 0.563 | 2.742 | 0.030 | 0.056 | 0.011 | 0.099 |
| Average | 0.364 | 1.861 | 0.040 | 0.043 | 0.012 | 0.067 |
| † Monocular method. ⋆ Scale-aligned with ground truth. | ||||||
IV-C Bird’s Eye View Prediction
is an efficient way of representing the environments in the top-down view. This representation provides a comprehensive spatial understanding, facilitates the fusion of various sensor modalities, and is easily adaptable for downstream tasks such as object tracking and planning. Networks predicting semantic maps in structured urban environments [36, 37] and predicting elevation and traversability maps in unstructured environments [38, 39, 40] have gained popularity in recent years. TartanGround, with its diversity and scale, provides an ideal platform to advance these approaches.
IV-D Neural Scene Representation
The photorealism of our dataset makes it well-suited for advancing research in neural scene representation techniques, such as Gaussian splatting [41] and NeRFs [42], as well as neural SLAM methods [43, 44, 45]. Its large-scale scenarios, with dynamic lighting and adverse weather conditions, offer challenging test cases for novel view synthesis and robust scene reconstruction. As shown in Fig. 7, novel view synthesis using Gaussian splatting trained on trajectories from the Rome and CoalMine environments demonstrates the potential of TartanGround for high-fidelity neural rendering.
IV-E Navigation
The dataset can be used for various advanced navigation techniques in ground robotics. Recent research has demonstrated the potential of imitation learning and diffusion-based approaches for robust navigation in complex environments. Models such as NoMaD [46], iPlanner [47], and ViPlanner [48] rely on diverse and large-scale datasets for training and evaluation. The TartanGround dataset, with its variety of environments, provides an ideal platform to develop and test these navigation models, enabling the creation of more robust and generalizable navigation systems for ground robots.
V Conclusion and Future Work
In this paper, we introduce TartanGround, a large-scale, multi-modal dataset designed to advance perception and navigation for ground robots in diverse environments. Our evaluations reveal the limitations of existing methods when applied to complex, unstructured environments, emphasizing the need for more robust and generalizable models. By providing comprehensive sensory data including multiple RGB stereo images, depth, semantic labels, LiDAR point clouds, semantic occupancy maps, and ground truth pose, TartanGround provides an ideal platform for training and benchmarking novel methods in occupancy prediction, SLAM, neural scene representation, visual navigation and more. In future, we aim to curate and release task-specific benchmarks utilizing the TartanGround dataset.
Acknowledgements
The authors would like to thank Jannick Schröer for generating the Gaussian splatting renderings.

### v2

@@
 Data is collected using an integrated automatic pipeline, which generates trajectories mimicking the motion patterns of various ground robot platforms, including wheeled and legged robots.
-We collect 910 trajectories across 70 environments, resulting in 1.5 million samples.
+We collect 878 trajectories across 63 environments, resulting in 1.44 million samples.
 Evaluations on occupancy prediction and SLAM tasks reveal that state-of-the-art methods trained on existing datasets struggle to generalize across diverse scenes.
 TartanGround can serve as a testbed for training and evaluation of a broad range of learning-based tasks, including occupancy prediction, SLAM, neural scene representation, perception-based navigation, and more, enabling advancements in robotic perception and autonomy towards achieving robust models generalizable to more diverse scenarios.
-The dataset and codebase for data collection will be made publicly available upon acceptance.
-Webpage: https://tartanair.org/tartanground
+The dataset and codebase are available on the webpage: https://tartanair.org/tartanground
 I Introduction
@@
 To address the need for comprehensive data resources, we introduce TartanGround, a large-scale simulation dataset designed to support perception and navigation tasks for ground robots across diverse environments.
-TartanGround includes data from over 70 diverse, challenging photorealistic environments, offering multi-modal sensor data such as multiple stereo RGB-D images covering a 360°, semantic labels, LiDAR point clouds, semantic occupancy maps, and ground truth poses.
+TartanGround includes data from over 63 diverse, challenging photorealistic environments, offering multi-modal sensor data such as multiple stereo RGB-D images covering a 360°, semantic labels, LiDAR point clouds, semantic occupancy maps, and ground truth poses.
 We show that the existing occupancy prediction methods, solely trained on data from the autonomous driving domain, do not generalize to environments of mobile robot operations such as forests.
 Moreover, we evaluate various algorithms and find that they struggle in challenging scenarios of low visibility and heavy occlusions.
-To summarize, the main contributions of this paper are: (1) A large dataset with over 1.5 million samples collected across diverse environments with precise ground truth labels mimicking different ground robot motion patterns, including wheeled (omnidirectional and differential-drive), and legged robots (quadrupedal), (2) an automatic data collection pipeline and (3) evaluation of two tasks, occupancy prediction and highlighting the limitations of models trained on existing datasets.
+To summarize, the main contributions of this paper are: (1) A large dataset with over 1.44 million samples collected across diverse environments with precise ground truth labels mimicking different ground robot motion patterns, including wheeled (omnidirectional and differential-drive), and legged robots (quadrupedal), (2) an automatic data collection pipeline and (3) evaluation of two tasks, occupancy prediction and highlighting the limitations of models trained on existing datasets.
 II Related Work
@@
 TartanGround is the only large-scale dataset covering diverse environments and having realistic ground robot motions.
-| Dataset | Env | Motion Pattern | Stereo | Depth | LiDAR | Semantics | Occupancy | Pose | Multi-cam | Seq Num | Samples |
-| Semantic KITTI [21] | Urban | Car | ✔ | ✘ | ✔ | ✔ | ✔ | ✔ | ✘ | 22 | 43 K |
-| Occ3D-Waymo [2, 22] | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 5 | 1000 | 200 K |
-| Occ3D-nuScenes [4, 22] | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 6 | 1000 | 40 K |
-| RELLIS-3D [5] | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 6.2 K |
-| RUGD [6] | Off-road | Car | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | 18 | 7.5 K |
-| TartanDrive [8] | Off-road | Car | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | 630 | 184 K |
-| WildScenes [10] | Off-road | Hand | ✘ | ✘ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 9.6 K |
-| WildOcc [25] | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✘ | 5 | 10 K |
-| GOOSE [7] | Off-road | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 356 | 10 K |
-| Scannet [11] | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✔ | ✔ | ✘ | 1500 | 2.5 M |
-| SUN RGBD [23] | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | - | 10 K |
-| GrandTour [27] | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✘ | ✘ | ✔ | 3 | 100 | 100 K |
-| TartanAir [14] | Mix | Random | ✔ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | 1037 | 400 k |
-| TartanAir-V2 [26] | Mix | Random | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 1110 | 1.4 M |
-| TartanGround | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 910 | 1.5 M |
+| Dataset | Type | Env | Motion Pattern | Stereo | Depth | LiDAR | Semantics | Occupancy | Pose | Multi-cam | Seq Num | Samples |
+| Semantic KITTI [21] | Real | Urban | Car | ✔ | ✘ | ✔ | ✔ | ✔ | ✔ | ✘ | 22 | 43 K |
+| Occ3D-Waymo [2, 22] | Real | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 5 | 1000 | 200 K |
+| Occ3D-nuScenes [4, 22] | Real | Urban | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 6 | 1000 | 40 K |
+| RELLIS-3D [5] | Real | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 6.2 K |
+| RUGD [6] | Real | Off-road | Car | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | 18 | 7.5 K |
+| TartanDrive [8] | Real | Off-road | Car | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | 630 | 184 K |
+| WildScenes [10] | Real | Off-road | Hand | ✘ | ✘ | ✔ | ✔ | ✘ | ✔ | ✘ | 5 | 9.6 K |
+| WildOcc [25] | Real | Off-road | Car | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | ✘ | 5 | 10 K |
+| SynPhoRest [27] | Syn. | Off-road | Random | ✘ | ✔ | ✔ | ✔ | ✘ | ✘ | ✘ | X | 3.1 K |
+| GOOSE [7] | Real | Off-road | Car | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 356 | 10 K |
+| GOOSE-Ex [28] | Real | Off-road | Tracked, Legged | ✘ | ✘ | ✔ | ✔ | ✔ | ✔ | 4 | 100 | 5 K |
+| Scannet [11] | Real | Indoor | Hand | ✘ | ✔ | ✘ | ✔ | ✔ | ✔ | ✘ | 1500 | 2.5 M |
+| GrandTour [29] | Real | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✘ | ✘ | ✔ | 3 | 100 | 100 K |
+| TartanAir [14] | Syn. | Mix | Random | ✔ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | 1037 | 400 k |
+| TartanAir-V2 [26] | Syn. | Mix | Random | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 1110 | 1.4 M |
+| TartanGround | Syn. | Mix | Legged, Wheeled | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 6 | 878 | 1.4 M |
 III The Dataset
 III-A Features
-TartanGround consists of 70 realistic simulation environments from TartanAir-V2 [26], covering diverse scenarios from structured urban to large-scale unstructured outdoor environments (Fig.
+TartanGround consists of 63 realistic simulation environments from TartanAir-V2 [26], covering diverse scenarios from structured urban to large-scale unstructured outdoor environments (Fig.
 2).
-The environments have been developed with Unreal Engine, and the data is collected using the AirSim [28] plugin.
+The environments have been developed with Unreal Engine 4, and the data is collected using the AirSim [30] plugin.
 This setup allows for rendering photorealistic scenes with high fidelity and makes it possible to have dynamic lighting, adverse weather effects, dynamic objects, and seasonal changes.
@@
 During post-processing, we generate additional ground truth data, including IMU data, optical flow, stereo disparity, LiDAR point clouds, and semantic occupancy maps.
-In total, we collect 910 trajectories (441 omni-wheeled, 226 diff-wheeled, and 243 legged), each with 600 to 8000 samples, resulting in a total dataset size of approximately 15 TB, 1.5 million samples, and 18 million RGB images.
+In total, we collect 878 trajectories (440 omni-wheeled, 198 diff-wheeled, and 240 legged), each with 600 to 8000 samples, resulting in a total dataset size of approximately 15 TB, 1.44 million samples, and 17.3 million RGB images.
 This large-scale, multi-modal dataset is designed to support a wide range of robotic perception and navigation tasks, benefiting the community in establishing new benchmarks for generalizable learning-based methods.
@@
 In this section, we show a small-scale evaluation of three state-of-the-art VO/SLAM algorithms on four challenging trajectories from four different environments.
+We use the front camera for these experiments.
 The trajectory from the Forest environment contains occlusion from tall grass and bushes, which are commonly presented in real-world off-road scenes.
@@
 IV-D Neural Scene Representation
-The photorealism of our dataset makes it well-suited for advancing research in neural scene representation techniques, such as Gaussian splatting [41] and NeRFs [42], as well as neural SLAM methods [43, 44, 45].
+The photorealism of our dataset makes it well-suited for advancing research in neural scene representation techniques, such as Gaussian splatting [43] and NeRFs [44], as well as neural SLAM methods [45, 46, 47, 48].
 Its large-scale scenarios, with dynamic lighting and adverse weather conditions, offer challenging test cases for novel view synthesis and robust scene reconstruction.
@@
 Acknowledgements
+This work was supported by the Luxembourg National Research Fund (Ref.
+18990533) and the Swiss National Science Foundation (Ref.
+200021E_229503).
+This work used Bridges-2 at PSC through allocation cis220039p from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program which is supported by NSF grants 2138259, 2138286, 2138307, 2137603, and 2138296.
 The authors would like to thank Jannick Schröer for generating the Gaussian splatting renderings.